# Multi-Modal Artificial Intelligence of Embryo Grading and Pregnancy Prediction in Assisted Reproductive Technology: A Review

### v1

@@ IV-B Pregnancy prediction
Although deep learning has made great progress in the task of embryo grading and has surpassed that of humans, the purpose of embryo grading is for pregnancy prediction. The role of embryo grading in pregnancy prediction is indirect, so current research tends to directly predict pregnancy outcomes. Pregnancy prediction is to predict the success of pregnancy after embryo transfer by analyzing the parents’ clinical data, embryo characteristics data, and laboratory test data. Current AI-based pregnancy prediction tasks can be divided into image-based, table-based and multi-modal based tasks according to different data types. Image-based task can be further divided into static image-based and temporal image-based. One point to be stated is that in this section we classify all three types of pregnancy, hCG biochemical pregnancy, fetal heart clinical pregnancy, and live birth, as pregnancies. As shown in Fig. 3, in the ”Task” column, Figure (e) represents the negative and positive embryos for pregnancy prediction based on the third day or the fifth day.
IV-B1 Pregnancy prediction based on static image
@@
The earlier machine learning method for pregnancy prediction based on static embryo images can be traced back to 2014. Xu et al. [72] extracted the features of static embryo images to 256 dimensions by Local Binary Pattern (LBP) [94]. Then Support Vector Machine (SVM) [95] was used to classify the extracted features, and the comprehensive performance of the test data in six institutions was ACC=64.71. In 2021, Geller et al. [73] collected 361 blastocyst stage embryo images and used the Inception V1 network to predict pregnancy. The final Area Under receiver operating characteristic Curve (AUC) was 65.70. Sawada et al. [74] fused the static embryo image information extracted by CNN branch and attention mechanism branch for pregnancy prediction, and the AUC reached 64.20.
IV-B2 Pregnancy prediction based on temporal image
@@
Previous image-based pregnancy prediction studies generally use images of the last day of embryo development to predict pregnancy, which ignores the timing of embryo development. Therefore, pregnancy prediction studies based on temporal embryo images have been gradually developed. Berntsen et al. [75] used I3D network [96] (a 3DCNN method) to obtain the features of 128 sequences of embryo development images, and then reduced the feature dimension by pooling operation. Finally, bidirectional LSTM [97] was used to connect temporal features for the final pregnancy prediction.
@@
Unlike image analysis, video analysis requires spatial-temporal data analysis, but many methods for processing spatial-temporal data such as 3DCNN are not suitable for medical applications due to high computational complexity. Therefore, Abbasi et al. [76] divided the pregnancy prediction task of temporal embryo image into three stages. In the first stage, CNN was used to extract individual image features on day 3 and day 5, and in the second stage, the temporal image features were fused through the attention mechanism in Transformer. In the third stage, the results of day 3 and day 5 time-series images were integrated by voting method to make the final pregnancy prediction. When processing TLM images, many frames of images will be recorded for each embryo, but these images have a lot of redundant information because the changes of embryos are small in a short time. Borna et al. [77] only used the three key development images of the first, second and third days of embryo development for pregnancy prediction. The ACC of the model reached 75.00, far higher than 60.32 of professional embryologists. Also using keyframe temporal images, Ouyang et al. [78] proposed a spatial-temporal position encoding, making the Transformer more suitable for fusing temporal images.
In order to understand the process of embryo development more comprehensively, Mapstone et al. [79] carried out the embryo grading task and pregnancy prediction task at the same time, and finally the ACC of the embryo 5-grading task was 88.00, and the AUC of pregnancy prediction was 68.00.
IV-B3 Pregnancy prediction based on tabular data
@@
In clinical practice, reproductive doctors need to predict pregnancy outcomes not only based on embryo images, but also based on tabular data such as parental fertility indicators. Tabular data is usually structured and easily quantified, and does not require the collection of large amounts of data like image processing tasks. Liu et al. [80] collected 401 cases of data with 19 fertility indicators, such as female age, female BMI, and EndoMetrial Thickness (EMT), and used SVM, Logistic Regression (LR) [98], Decision Tree (DT) [99], and Random Forest (RF) [100] to predict pregnancy. The AUC of the best RF method was 61.30, and it was found through statistics that: Among patients with good-quality embryos, the clinical pregnancy success rate was about 70 when EMT thickness was greater than 9.6 mm, but 50 when EMT thickness was no greater than 9.6 mm. Mehta et al. [81] collected data of 333 cases with 11 indicators such as female age, Antral Follicle Count (AFC), and Follicle-Stimulating Hormone (FSH), and observed that the ACC of pregnancy prediction of Adaboost [101] was 97.50, better than all other machine learning models, and found statistically that: The propensity to clinical pregnancy is negative if the woman is older than 36 years, and the likelihood of clinical pregnancy decreases substantially if the woman is older than 40 years; Furthermore, the propensity for clinical pregnancy was positively correlated with the number of embryos transferred in the same IVF cycle.
@@
Similarly, through statistical related studies [82][83][84][85][86][87], it can be concluded that the characteristics affecting pregnancy prediction mainly include female age, Anti-Mullerian Hormone (AMH), EMT, semen volume, oocytes number, AFC, female BMI and so on. In addition to using traditional machine learning methods for pregnancy prediction, Cao et al. [88] converted 123 relevant data into natural language and fine-tuned the large language model GPT-4 with the converted natural language, resulting in a final AUC of 87.00. Ouyang et al. [78] converted the tabular data into tokens and input them into Transformer for pregnancy prediction with AUC of 71.30.
IV-B4 Pregnancy prediction based on multi-modal data
@@
With the progress of AI and the improvement of medical data collection technology, pregnancy prediction is more likely to use multi-modal data such as images and tables. This trend is mainly due to the potential of multi-modal data fusion, which integrates the advantages of different types of data to provide more comprehensive prediction results, thereby improving the clinical application value of pregnancy prediction. Compared with image data, tabular data can directly express high-level semantics information. Therefore, some works [89][90] extract image information into low-dimensional data through CNN, and then input it into the fusion model combined with tabular data directly. The quality of the embryo at the blastocyst stage is mainly reflected from two aspects, ICM and TE. Based on this prior knowledge, Kim et al. [91] first binarized and segmented the ICM and TE of the embryo through the traditional image segmentation method, and then added the segmentation map to the original image to obtain the embryo enhancement image. Finally, the features extracted from the enhancement image were combined with age features in the pregnancy prediction task. Similarly, Liu et al. [92] collected 17580 multi-modal data for pregnancy prediction, in which the image data contained two images focusing on ICM and TE respectively, and the tabular data contained 16 parental fertility indicators.
@@
Considering that the image modality data used in the previous multi-modal pregnancy study are all static, Kim et al. [93] collected the time series embryos image data from day 1 to day 5 and health spreadsheet data. When processing the image data, it was multi-stage. Firstly, each morphology of the embryo was segmented through the segmentation model. Then the segmentation images were input into the neural network to predict five key indicators: blastocyst size, blastocyst level, cell boundary, cell number and developmental stage. After processing, the original time series image, spreadsheet data, morphological segmentation image and five key indicators were input into Transformer for final pregnancy prediction. In addition, Ouyang et al. [78] are the first to integrate images of the first 3 days of embryo development and fertility table data for pregnancy prediction, and propose an effective decoupling fusion network to fuse multi-modal data.
In terms of the performance of pregnancy prediction task, models based on tabular data generally significantly outperform models using only images. This may be due to the closer temporal distance and stronger correlation between tabular data and pregnancy outcomes, thus enabling the model to capture more direct and critical predictive features. In addition, tabular data are often extracted by clinical experts with highly correlated features of pregnancy outcomes, which has strong prior value and further improves the performance of the model. It is worth noting that even in the multi-modal modeling scenario, the tabular model still generally outperforms the multi-modal methods that fuse images and tables. This phenomenon may not result from the ineffectiveness of multi-modal fusion itself, but more reflects the realistic challenges existing in current multi-modal research. For example, existing work is mostly based on private datasets, and there are significant differences among different institutions in data collection processes, feature encoding methods, and image device resolutions, which leads to the lack of comparability of model performance under different data sources. The details of these methods are shown in the TABLE III.
| Dataset | Year | Modality | Description | Download link |
| A time-lapse embryo dataset for morphokinetic parameter prediction.[102] | 2022 | Static image | The dataset consists of 704 time-lapse videos of developing embryos with 7 focal length, covering day 2-6 of embryo development, for a total of 2.4 million images. Annotations are graded to 16 different developmental stages. | https://zenodo.org/records/6390798 |
| Embryologist agreement when assessing blastocyst implantation probability: is data-driven prediction the solution to embryo assessment subjectivity?[44] | 2022 | Temporal image | The dataset consists of 143 time-lapse videos of embryo development, including day 1-5 of embryo development. Annotations are 36 score ratings annotated by embryologists, which can be further divided into three grades: good, medium, and bad. | https://dataverse.harvard.edu/dataset.xhtml?persistentId=doi:10.7910/DVN/1Z4HLC |
| Adaptive adversarial neural networks for the analysis of lossy and domain-shifted datasets of medical images.[103] | 2022 | Static image | The dataset consists of 2440 images, including day 5 images of embryo development. The annotation is the five ratings 1-5, where ratings 1 and 2 are for no blastocysts developed and ratings 3,4 and 5 are for blastocysts developed. | https://osf.io/3kc2d/ |
| Embryo classification based on microscopic images.[104] | 2023 | Static image | The dataset consists of 1020 images, including images on day 3 and day 5 of embryo development. The annotations are good and bad ratings. | https://www.kaggle.com/competitions/world-championship-2023-embryo-classification/overview |
| Merging synthetic and real embryo data for advanced AI predictions.[105] | 2025 | Static image | The dataset consists of 5500 images generated by AI, including images from day 1-5 of embryo development. The annotations are the five grades 2-cell, 4-cell, 8-cell, morula, and blastocyst. | https://zenodo.org/records/14253170, https://huggingface.co/datasets/deepsynthbody/synembryo_latentdiffusion, https://huggingface.co/datasets/deepsynthbody/synembryo_stylegan |
| Automatic Identification of Human Blastocyst Components via Texture.[56] | 2017 | Static image | The dataset consists of 211 images, including day 5 images of embryo development. Annotations are segmented labels for different regions of the blastocyst, namely ICM, TE, BC, and ZP. | https://vault.sfu.ca/index.php/s/066vGJfviJMYuP6/authenticate |
| Cleavage-stage embryo segmentation using SAM-based dual branch pipeline: development and evaluation with the CleavageEmbryo dataset.[69] | 2024 | Static image | The dataset consists of 1548 images, including images from day 1 to 3 of embryo development. Annotation is the COCO object detection format, which detects objects separately for embryos, fragments, and background. | https://www.kaggle.com/datasets/austin012/cleavageembryo-dataset/data |
| Neural networks pipeline for quality management in IVF laboratory.[106] | 2024 | Table | The dataset consists of 11 1500 treatment cycles and includes 19 key tabular data. Label is the result of fetal heart clinical pregnancy. | https://github.com/embryossa/KAN-in-IVF |
| An annotated human blastocyst dataset to benchmark deep learning architectures for in vitro fertilization.[107] | 2023 | Static image and Table | The dataset consists of 2344 images from 837 cases, including images on day 4 or 5 of embryo development. These 2344 images were annotated with Expansion 5-grading, ICM 3-grading, and TE 3-grading. In addition, the 837 cases included six tabular features for hCG biochemical pregnancy, fetal heart clinical pregnancy, and live birth annotation. | https://figshare.com/articles/figure/Blastocyst_dataset_zip/20123153/3?file=39348899 |
IV-C Explainability
@@
In ART, AI models are widely applied in embryo quality assessment, pregnancy success rate prediction and other key diagnosis and treatment tasks. Although these models show significant advantages in performance, their ”black box” feature makes explainability analysis be a prerequisite for clinical deployment. On the one hand, ART involves highly sensitive issues such as bioethics, patients’ right to informed consent, and the attribution of medical responsibility. The decision-making basis of the model must be transparent to doctors and patients. On the other hand, explanatory analysis not only helps enhance doctors’ trust in the prediction results of AI, but also helps identify potential medical-related features.
@@
SHapley Additive exPlanations (SHAP) [109] can quantitatively analyze the predictive contributions of input variables, especially suitable for pregnancy prediction models constructed based on clinical tabular data, such as evaluating the positive and negative effects and weights of factors such as age, hormone levels, and embryo scores on the output probability [78][89].
Furthermore, for image-based model explainability, Gradient-weighted Class Activation Mapping (Grad-CAM) highlights the regions that the model is concerned about in the image by generating heat maps. Grad-CAM helps researchers to determine whether the AI has focused on clinically significant structures (such as ICM or TE) in embryonic images, thereby evaluating the rationality and usability of its decision-making process [35][110]. Overall, these interpretable technologies provide crucial support for AI applications in the ART field. They not only enhance the transparency of predictive models but also offer important clues for model optimization and clinical integration.
IV-D Public dataset

### v2

@@ 4.2 Pregnancy prediction
Although deep learning has made great progress in the task of embryo grading and has surpassed that of humans, the purpose of embryo grading is for pregnancy prediction. The role of embryo grading in pregnancy prediction is indirect, so current research tends to directly predict pregnancy outcomes. Pregnancy prediction is to predict the success of pregnancy after embryo transfer by analyzing the parents’ clinical data, embryo characteristics data, and laboratory test data. Current AI-based pregnancy prediction tasks can be divided into image-based, table-based and multi-modal based tasks according to different data types. Image-based task can be further divided into static image-based and temporal image-based. One point to be stated is that in this section we classify all three types of pregnancy, hCG biochemical pregnancy, fetal heart clinical pregnancy, and live birth, as pregnancies. As shown in Fig. 3, in the ”Task” column, Fig.3 (e) represents the negative and positive embryos for pregnancy prediction based on the third day or the fifth day.
4.2.1 Pregnancy prediction based on static image
@@
The earlier machine learning method for pregnancy prediction based on static embryo images could be traced back to 2014. Xu et al. [101] extracted the features of static embryo images to 256 dimensions by Local Binary Pattern (LBP) [135]. Then Support Vector Machine (SVM) [136] was used to classify the extracted features, and the comprehensive performance of the test data in six institutions was ACC=64.71. Miyagi et al. [102] collected 180 blastocyst images and used a Logistic Regression (LR) model combined with five-fold cross-validation to predict pregnancy outcomes, achieving an Area Under receiver operating characteristic Curve (AUC) of 65.00. Geller et al. [103] collected 361 blastocyst stage embryo images and used the InceptionV1 network to predict pregnancy. The final AUC was 65.70. Sawada et al. [104] fused the static embryo image information extracted by CNN branch and attention mechanism branch for pregnancy prediction, and the AUC reached 64.20. Rad et al. [105] employed a two-stream network to extract the features of the blastocyst segmentation image and the original image respectively, and then fused them for pregnancy prediction. The final accuracy rate was 70.90.
4.2.2 Pregnancy prediction based on temporal image
@@
Previous image-based pregnancy prediction studies generally used images of the last day of embryo development to predict pregnancy, which ignored the timing of embryo development. Therefore, pregnancy prediction studies based on temporal embryo images had been gradually developed. Berntsen et al. [106] used I3D network [137] (a 3DCNN method) to obtain the features of 128 sequences of embryo development images, and then reduced the feature dimension by pooling operation. Finally, bidirectional LSTM [138] was used to connect temporal features for the final pregnancy prediction.
@@
Unlike image analysis, video analysis requires spatial-temporal data analysis, but many methods for processing spatial-temporal data such as 3DCNN are not suitable for medical applications due to high computational complexity. Therefore, Abbasi et al. [33] divided the pregnancy prediction task of temporal embryo image into three stages. In the first stage, CNN was used to extract individual image features on day 3 and day 5, and in the second stage, the temporal image features were fused through the attention mechanism in Transformer. In the third stage, the results of day 3 and day 5 time-series images were integrated by voting method to make the final pregnancy prediction. When processing TLM images, many frames of images would be recorded for each embryo, but these images had a lot of redundant information because the changes of embryos were small in a short time. Borna et al. [107] only used the three key development images of the first, second and third days of embryo development for pregnancy prediction. The ACC of the model reached 75.00, far higher than 60.32 of professional embryologists. Also using keyframe temporal images, Ouyang et al. [108] proposed a spatial-temporal position encoding, making the Transformer more suitable for fusing temporal images. In order to understand the process of embryo development more comprehensively, Mapstone et al. [109] carried out the embryo grading task and pregnancy prediction task at the same time, and finally the ACC of the embryo 5-grading task was 88.00, and the AUC of pregnancy prediction was 68.00.
Boucret et al. [110] first extracted the video features based on the self-supervised learning method, and then used XGBoost. The final AUC was 64.00. In clinical practice, the label of pregnancy negative was less dependent on the quality of the embryo, which would cause the result of model overfitting when training. To solve this problem, Nagaya et al. [111] used the method of unlabeled learning to change the label of pregnancy negative to unlabeled, which enhanced the regularization of the model and improved the performance of the model.
4.2.3 Pregnancy prediction based on tabular data
@@
In clinical practice, reproductive doctors needed to predict pregnancy outcomes not only based on embryo images, but also based on tabular data such as parental fertility indicators. Tabular data is usually structured and easily quantified, and did not require the collection of large amounts of data like image processing tasks. Liu et al. [112] collected 401 cases of data with 19 fertility indicators, such as female age, female BMI, and EndoMetrial Thickness (EMT), and used SVM, LR [139], Decision Tree (DT) [140], and Random Forest (RF) [141] to predict pregnancy. The AUC of the best RF method was 61.30, and it was found through statistics that: Among patients with good-quality embryos, the clinical pregnancy success rate was about 70 when EMT thickness was greater than 9.6 mm, but 50 when EMT thickness was no greater than 9.6 mm. Mehta et al. [113] collected data of 333 cases with 11 indicators such as female age, Antral Follicle Count (AFC), and Follicle-Stimulating Hormone (FSH), and observed that the ACC of pregnancy prediction of Adaboost [142] was 97.50, better than all other machine learning models, and found statistically that: The propensity to clinical pregnancy was negative if the woman was older than 36 years, and the likelihood of clinical pregnancy decreased substantially if the woman was older than 40 years; Furthermore, the propensity for clinical pregnancy was positively correlated with the number of embryos transferred in the same IVF cycle.
@@
Similarly, through statistical related studies [114][115][116][117][118][119], it can be concluded that the characteristics affecting pregnancy prediction mainly include female age, Anti-Mullerian Hormone (AMH), EMT, semen volume, oocytes number, AFC, female BMI and so on. In addition to using traditional machine learning methods for pregnancy prediction, Cao et al. [120] converted 123 relevant data into natural language and fine-tuned the large language model GPT-4 with the converted natural language, resulting in a final AUC of 87.00. Ouyang et al. [108] and Borji [121] et al. converted the tabular data into tokens and input them into Transformer for pregnancy prediction. Liu et al. [122] transformed one-dimensional tabular data into two-dimensional data through replication operations and then extracted features through CNN for pregnancy prediction.
The core characteristics and implementation details of the aforementioned studies are summarized in TABLE 4. To broaden the scope and enhance the comprehensiveness of this review, we further incorporate a number of related works that share similar technical methodologies or target comparable application tasks [121][123][124][125][126][127].
4.2.4 Pregnancy prediction based on multi-modal data
@@
With the progress of AI and the improvement of medical data collection technology, pregnancy prediction was more likely to use multi-modal data such as images and tables. This trend was mainly due to the potential of multi-modal data fusion, which integrated the advantages of different types of data to provide more comprehensive prediction results, thereby improving the clinical application value of pregnancy prediction. Compared with image data, tabular data could directly express high-level semantics information. Therefore, some works [128][129][130] extract image information into low-dimensional data through CNN, and then input it into the fusion model combined with tabular data directly. The quality of the embryo at the blastocyst stage is mainly reflected from two aspects, ICM and TE. Based on this prior knowledge, Kim et al. [131] first binarized and segmented the ICM and TE of the embryo through the traditional image segmentation method, and then added the segmentation map to the original image to obtain the embryo enhancement image. Finally, the features extracted from the enhancement image were combined with age features in the pregnancy prediction task. Similarly, Liu et al. [132] collected 17580 multi-modal data for pregnancy prediction, in which the image data contained two images focusing on ICM and TE respectively, and the tabular data contained 16 parental fertility indicators.
@@
Considering that the image modality data used in the previous multi-modal pregnancy study were all static, Kim et al. [133] collected the time series embryos image data from day 1 to day 5 and health spreadsheet data. When processing the image data, it was multi-stage. Firstly, each morphology of the embryo was segmented through the segmentation model. Then the segmentation images were input into the neural network to predict five key indicators: blastocyst size, blastocyst level, cell boundary, cell number and developmental stage. After processing, the original time series images, spreadsheet data, morphological segmentation images and five key indicators were input into Transformer for final pregnancy prediction. In addition, Ouyang et al. [108] were the first to integrate images of the first 3 days of embryo development and fertility table data for pregnancy prediction, and proposed an effective decoupling fusion network to fuse multi-modal data. Dual et al. used 3DResNet to extract video data and combined it with 31 tabular data for the final pregnancy assessment. The details of these methods are shown in the TABLE 4.
4.3 Summary of current works
Fig. 4(a) visually presents the distribution relationships among different periods (cleavage stage and blastocyst stage), different modalities (static images, temporal images, tables and multi-modalities), and different task types (embryo grading (EG), cell counting (CC), embryo segmentation (ES) and pregnancy prediction (PP)) in the form of a sunburst chart. It can be clearly seen from the figure that the number of studies related to embryo grading is significantly greater than that of pregnancy prediction tasks, and the research work at the blastocyst stage is far more than that at the cleavage stage. This distribution feature indicates that the research related to pregnancy prediction during the cleavage stage is relatively weak. Future studies may consider strengthening the exploration in these directions.
Fig. 4(b) delineates the technical methodologies AI in ART through a tree diagram. Synthesizing this visualization with the current studies reveals several prominent technical routes, which are primarily determined by the data modality. For static image analysis, the predominant approach involves feature extraction using foundational deep learning architectures, including CNNs, Transformers, and GNNs. In contrast, the analysis of temporal image sequences necessitates models capable of capturing dynamic information, such as RNNs, 3D CNNs, and Transformer-based architectures. For structured tabular data, conventional yet highly efficient machine learning algorithms are typically employed. In addition, some methods take a different approach, using language models or CNNs to handle tabular modal data.
Beyond these foundational models, several advanced strategies are consistently utilized to augment predictive performance. Ensemble learning and attention mechanisms have proven effective in significantly boosting model accuracy. Furthermore, to address the inherent complexity of ART tasks, researchers often decompose problems into manageable sub-tasks using strategies like incorporating prior medical knowledge, implementing two-stage models, or designing multi-network frameworks. Finally, emerging techniques such as self-supervised learning show considerable promise for enhancing model representational ability, particularly in research contexts constrained by limited data availability.
Fig. 4(c) presents the statistical distribution of performance for different tasks in the form of a box plot. Due to the differences in the evaluation criteria of embryo grading tasks, the performance variance is relatively large. However, the average ACC of embryo grading under both static images and temporal images exceeds 75, and it has significantly outperformed the subjective assessment level of embryologists [42][43][47].
At present, a variety of mature products have been put into clinical use, further promoting the standardization and intelligence of embryo grading in the field of assisted reproduction. STORK (https://github.com/ih-lab/STORK) is an AI model from Stanford University, trained on the Inception image recognition architecture. The model can identify the morphological characteristics of the embryo and output the transfer probability or classification results. Other commercial software scoring systems, such as KIDScore and iDAScore [143], are also widely used. They are fully AI-driven scoring systems developed by Vitrolife, outputting a continuous score (e.g., 1 to 10) for prioritization of embryos. Meanwhile, ERICA (Embryo Ranking Intelligent Classification Algorithm) [144] is also an embryo evaluation model based on deep learning, aiming to automatically score and rank static embryos in ART, which can be accessed on https://embryoranking.com/.
In addition, when conducting pregnancy prediction tasks based on tabular data, due to the significant differences in the clinical features included by different hospitals or institutions, coupled with the non-uniformity of testing equipment, data collection methods, and standardized processing procedures, the model performance shows a large variance as shown in Fig. 4(c). At present, there is no recognized standardized process for processing tabular data, which to some extent limits the generalization ability and clinical applicability of the model among different centers. And the pregnancy prediction model based on the table’s uni-modal model demonstrates superior predictive performance compared to the multi-modal model. This phenomenon may be related to the differences in the number of indicators used by different institutions: models based on tabular data use an average of 48 indicators, while multi-modal models use an average of only 15 indicators. The significant difference in the number of indicators may be one of the key factors leading to performance differences.
Although the tabular indicators adopted in different studies vary, in the task of pregnancy prediction, by comparing and analyzing the statistical performance of static images, temporal images and multi-modal methods as shown in Fig. 4(c), it can be clearly observed that the multi-modal method introducing tabular modality has significantly better performance than the method relying only on a single image modality (static or temporal). This result indicates that the clinical features contained in the tabular data are complementary to the image data. Fusing multi-modal data can effectively enhance the prediction accuracy and robustness of the model. The above findings are highly consistent with the results of [108][128][132]. To advance the clinical implementation of pregnancy prediction, it is imperative to establish standardized processing methodologies for tabular data modalities, integrate multi-modal data, and subsequently develop high-performance, highly reliable models leveraging large-scale datasets.
5 Explainability
@@
In ART, AI models are widely applied in embryo quality assessment, pregnancy success prediction and other key diagnosis and treatment tasks. Although these models show significant advantages in performance, their ”black box” feature makes explainability analysis be a prerequisite for clinical deployment. On the one hand, ART involves highly sensitive issues such as bioethics, patients’ right to informed consent, and the attribution of medical responsibility. The decision-making basis of the model must be transparent to doctors and patients. On the other hand, explanatory analysis not only helps enhance doctors’ trust in the prediction results of AI, but also helps identify potential medical-related features.
@@
To address the interpretability challenge inherent in image-based models, researchers have developed a range of visualization analysis techniques, as illustrated in Fig. 3. Among these, Gradient-weighted Class Activation Mapping (Grad-CAM) has emerged as a widely adopted technique. It operates by generating a heatmap that visually highlights the spatial regions upon which the model focuses its attention. This capability allows researchers to verify whether the AI system is attending to clinically significant structures within the embryo image, such as the inner cell mass (ICM) or the trophectoderm (TE), thereby enabling a critical evaluation of the rationale and clinical validity of its decision-making process [45][146]. Beyond Grad-CAM, other prominent methods offering similar functionalities include Local Interpretable Model-agnostic Explanations (LIME) [147] and SHapley Additive exPlanations (SHAP) [148] . These approaches further bolster the transparency and trustworthiness of model decisions by employing local approximation or feature attribution analysis, respectively [149].
Besides image models, SHAP and LIME are also suitable for interpretability analysis of tabular data. These two methods can quantitatively evaluate the contribution of each input variable to the model prediction results, and clarify the positive or negative influence of key factors such as patient age, hormone level, and embryo score on the output probability and their weight distribution [108][128]. Therefore, SHAP and LIME show significant advantages in processing multi-modal data, which can effectively explain the feature interactions from different modalities, and provide transparent and reliable explanation support for the decision-making process of complex models.
| Dataset | Year | Modality | Description |
| [150] | 2022 | Static image | The dataset consists of 704 time-lapse videos of developing embryos with 7 focal length, covering day 2-6 of embryo development, for a total of 2.4 million images. Annotations are graded to 16 different developmental stages. https://zenodo.org/records/6390798 |
| [63] | 2022 | Temporal image | The dataset consists of 143 time-lapse videos of embryo development, including day 1-5 of embryo development. Annotations are 36 score ratings annotated by embryologists, which can be further divided into three grades: good, medium, and bad. https://dataverse.harvard.edu/dataset.xhtml?persistentId=doi:10.7910/DVN/1Z4HLC |
| [151] | 2022 | Static image | The dataset consists of 2440 images, including day 5 images of embryo development. The annotation is the five ratings 1-5, where ratings 1 and 2 are for no blastocysts developed and ratings 3,4 and 5 are for blastocysts developed. https://osf.io/3kc2d/ |
| [152] | 2023 | Static image | The dataset consists of 1020 images, including images on day 3 and day 5 of embryo development. The annotations are good and bad ratings. https://kaggle.com/competitions/world-championship-2023-embryo-classification |
| [153] | 2025 | Static image | The dataset consists of 5500 images generated by AI, including images from day 1-5 of embryo development. The annotations are the five grades 2-cell, 4-cell, 8-cell, morula, and blastocyst. https://zenodo.org/records/14253170, https://huggingface.co/datasets/deepsynthbody/synembryo_latentdiffusion, https://huggingface.co/datasets/deepsynthbody/synembryo_stylegan |
| [83] | 2017 | Static image | The dataset consists of 211 images, including day 5 images of embryo development. Annotations are segmented labels for different regions of the blastocyst, namely ICM, TE, BC, and ZP. https://vault.sfu.ca/index.php/s/066vGJfviJMYuP6/authenticate |
| [97] | 2024 | Static image | The dataset consists of 1548 images, including images from day 1 to 3 of embryo development. Annotation is the COCO object detection format, which detects objects separately for embryos, fragments, and background. https://www.kaggle.com/datasets/austin012/cleavageembryo-dataset/data |
| [154] | 2024 | Table | The dataset consists of 8732 treatment cycles and includes 19 key tabular data. Label is the result of fetal heart clinical pregnancy. https://github.com/embryossa/KAN-in-IVF |
| [155] | 2023 | Static image and Table | The dataset consists of 2344 images from 837 cases, including images on day 4 or 5 of embryo development. These 2344 images were annotated with Expansion 5-grading, ICM 3-grading, and TE 3-grading. In addition, the 837 cases included six tabular features for hCG biochemical pregnancy, fetal heart clinical pregnancy, and live birth annotation. https://figshare.com/articles/figure/Blastocyst_dataset_zip/20123153/3?file=39348899 |
6 Public Dataset